OpenStudy (anonymous):

The region R bounded by the curve 4y=x² and the lines x=4, y=1. Find the volume of the solid formed when R is rotated completely about (a) the x-axis (b) the y-axis (c) the line y=1(d) the line x=4 the answers should be (a) 52pi/5 (b) 18pi (c) 76pi/15 (d)10pi/3

OpenStudy (kira_yamato):

For the first two, you need to remember two formulas: \[V_x=π\int\limits_{\inf \lim}^{\sup \lim}y^2 dx\] \[V_y=π\int\limits_{\inf \lim}^{\sup \lim}x^2 dy\] For the next two, you'll need to bring the curve to set the limits at the axes to apply the formulas by doing a corresponding translation of the curve

OpenStudy (anonymous):

ok thanks so can we go through step by step for the first two, i will try as much as i can and correct me if am wrong.

OpenStudy (anonymous):

4y=x^2 y=(x^2)/4 y^2=(x^4)/16 is this correct

OpenStudy (kira_yamato):

Yes

OpenStudy (anonymous):

so we integrate this and get x^5/80. how do we get the limits of integration?

OpenStudy (kira_yamato):

You get your limits from the question, though it'll be easier if you sketch out the graph

OpenStudy (anonymous):

i have got the sketch the point of intersection is what am struggling with

OpenStudy (anonymous):